This mass (m) on a string (of length L) is moving
Answer #1

Net force on m along vertical direction

Tcos(theta) = mg

T*cos(60) = mg

Tebsion in the string T = 2mg


From conservation of energy

KE = PE

(1/2)mv^2 = mgh

v^2 = 2gL(1-cos(theta)

v^2 = 2gL(1-cos(60))

v = (gL)^0.5